Find the greatest number of four digits which when divided by 10,11,15 and 22 leaves 3,4,8 and 15 as remainders respectively. (stepwise)
a)9907             b)9903
c)9983             d)none of these

Grade:12th pass

1 Answers

Harsh Patodia IIT Roorkee
9 years ago
10-3 =7
11-4= 7
15-8=7
22-15=7
As the difference between divisor and remainder is constant we subtract this from highest muliplt to get th desired result.
FInd LCM of 4 numbers and find the greatest multiple of that in 4 digits and subtract 7 from that.
LCM (10,11,15,22)= 330
Greatest Multiple of that in 4 digits = 9900
Subtract 7 from that you will get 9893.
So D is correct answer.